SQRT(7x+25) = x + 1
Found 2 solutions by tommyt3rd, stanbon:
Answer by tommyt3rd(5050)

SQRT(7x+25) = x + 1
7x+25=(x+1)^2
7x+25=x^2+2x+1

x^2-5x-24=0
(x-8)(x+3)=0
x=8 or x=-3 (extraneous)

x=8

:)



Answer by stanbon(75887)

SQRT(7x+25) = x + 1
------
Square both sides:
7x+25 = (x+1)^2
------
7x+25 = x^2+2x+1
------
x^2 -5x -24 = 0
Factor:
(x-8)(x+3) = 0
x = 8 or x = -3
------
x = -3 is extraneous
-----
x = 8 is the only solution
=================================
Cheers,
Stan H.
==========================